OpenStudy (anonymous):

Choose the correct simplification of the expression (3x)4

jimthompson5910 (jim_thompson5910):

Similar Example \[\large (2x)^5 = 2^5*x^5\] \[\large (2x)^5 = 32x^5\]

OpenStudy (anonymous):

I don't understand.

jimthompson5910 (jim_thompson5910):

effectively, in the example, we're raising each piece to the 5th power

jimthompson5910 (jim_thompson5910):

so (2x)^5 turns into 2^5 * x^5 we raise the 2 to the 5th power and the x to the 5th power

jimthompson5910 (jim_thompson5910):

2^5 is the same as 32 (2^5 = 2*2*2*2*2 = 32) so that's why 2^5 * x^5 turns into 32x^5

jimthompson5910 (jim_thompson5910):

this isn't the answer, but it's an example showing you how to do it

OpenStudy (anonymous):

Oh okay. I get how you did that.

jimthompson5910 (jim_thompson5910):

that's great, glad it's clicking now

OpenStudy (anonymous):

Thanks.

jimthompson5910 (jim_thompson5910):

yw